OpenStudy (jilltalks):

How do you solve this system? 1/3x +3/4y = 2 4/9x - 1/4y = 3

OpenStudy (sshayer):

multiply second eq by 3 and add

OpenStudy (jilltalks):

??????

OpenStudy (sshayer):

\[3*\frac{ 4 }{ 9 }x-3*\frac{ 1 }{ 4 }y=3*3\] \[\frac{ 4 }{ 3 }x-\frac{ 3 }{ 4 }y=9\] now add eq. (1)